A box contains 5 red and 7 green balls. How many ways can 2 balls be selected with the 2 balls the same color??

I am totally stumped???

Answer by stanbon(75887)

A box contains 5 red and 7 green balls. How many ways can 2 balls be selected with the 2 balls the same color??
-----------
Selecting 2 reds or selecting 2 greens mutually exclusive events.
Formula: P(A or B) = P(A)+P(B)
P(2red) = [number of ways to select 2 from 5]/[# of ways to select 2 from 12]
P(2green)=[# of ways to select 2 from 7]/[# of ways to select 2 from 12}
P(2 balls same) = P(2red)+P(2green)
=[5C2]/[12C2] + [7C2]/[12C2]
=(10+21)/66
=31/66
===========
Cheers,
Stan H.